Louisiana State University Health Sciences Center Administration & Finance
 

Friday, March 30, 2007

 

MEB:

  • MEB outage was performed on time and successfully. There was a problem getting the rooftop generator started, but it was able to be started by 7:30.
  • Facility Services personnel took advantage of the outage, and documented the emergency power source for the operating equipment for future reference.

Downtown Campus Landscaping Project:

  • Only One water meter remains to be installed. It is the one on the 500 block of South Roman Street .

Dental School: 

  • Schindler was on site continuing repairs on the elevators.
  •  Zimmer has finished removing ceiling grid on the 2nd floor of both buildings and has started on the 3rd floor. They are working this weekend

1542 Tulane Ave:

  • Ordes Electrical continues to install the 4" conduit to the vault.
  • Stratos Elevators continued working on the Elevator# 3 cab.
  • Mechanical Insulation is insulating chilled water piping in the basement.